# Impacts of aspect ratio on task accuracy in parallel coordinates

### v1

@@ VII Recommendations
Other methods should ideally be used in conjunction with a PCP [39], but this is a common task and PCPs used in exploratory analysis may present distorted findings. As such, we suggest the following: (i) for low sample sizes, a more balanced estimation of positive and negative correlations may be achieved using an AR greater than 1; (ii) small ARs are particularly poor for estimating +ve correlations and should ideally be avoided unless used with a linked view; (iii) where sample size is unknown, an AR of 2 is likely to provide the most balanced option for correlation estimation.
@@
Although value tracing is frequently used in combination with interaction methods such as brushing, users performing exploratory tasks may use rapid non-interactive visual scanning to focus investigations and later interaction. ARs of between 0.5 and 1 are likely to be most effective to enable accurate value tracing. As such, we suggest the following: (i) very small ($<$0.25) and very large (4) ARs strongly affect the performance of users in value tracing tasks and should be avoided, unless using brushing for fully-accurate value tracing, or enabling other interaction methods such as axis ordering and lensing techniques.
The above recommendations conflict if the PCP is intended (or is likely to be used) for both tasks. Primarily, a high AR providing balanced correlation estimation for low sample sizes will substantially reduce performance for value tracing. User information, interaction and linked views reduce this issue, although visualization designers will need to understand the primary use-case for the PCP.
VIII Future work
Similar to the optimisation algorithms proposed for line charts (for example [18] and [40]) or scatter plots (for example [15] or [41], we note that it may be possible to design an optimised custom spacing of axes. This would create an equally perceptually salient/identical discriminability of each z (i.e. varying the AR of the axes depending on a pre-calculated z), and could include adjustments for intended task.
Supplemental Materials
All supplemental materials are available on Figshare at 10.6084/m9.figshare.27013108.
Acknowledgments

### v2

@@ 7 Recommendations
Other methods should ideally be used in conjunction with a PCP [43], but this is a common task and PCPs used in exploratory analysis may present distorted findings. As such, we suggest the following: (i) for low sample sizes, a more balanced estimation of positive and negative correlations may be achieved using an aspect ratio greater than 1; (ii) small aspect ratios are particularly poor for estimating positive correlations and should ideally be avoided unless used with a linked view; (iii) where sample size is unknown, an aspect ratio of 2 is likely to provide the most balanced option for correlation estimation.
@@
Although value tracing is frequently used in combination with interaction methods such as brushing, users performing exploratory tasks may use rapid non-interactive visual scanning to focus investigations and later interaction. Aspect ratios of between 0.5 and 1 are likely to be most effective to enable accurate value tracing. As such, we suggest the following: (i) very small (<<0.25) and very large (4) aspect ratios strongly affect the performance of users in value tracing tasks and should be avoided, unless using brushing for fully-accurate value tracing, or enabling other interaction methods such as axis ordering and lensing techniques.
The above recommendations conflict if the PCP is intended (or is likely to be used) for both tasks. Primarily, a high aspect ratio providing balanced correlation estimation for low sample sizes will substantially reduce performance for value tracing. User information, interaction and linked views reduce this issue, although visualization designers will need to understand the primary use-case for the PCP.
8 Future work
Similar to the optimisation algorithms proposed for line charts (for example [19] and [44]) or scatter plots (for example [16] or [45], we note that it may be possible to design an optimised custom spacing of axes. This would create an equally perceptually salient/identical discriminability of each z (i.e. varying the aspect ratio of the axes depending on a pre-calculated z), and could include adjustments for intended task.
The PCP is not necessarily ideally suited to correlation estimation tasks and supplementary visualizations or methods exist to mitigate some of the issues in correlation perception (such as ’angle uniform parallel coordinates’ [46] or ’indexed points parallel coordinates’ [47]). However, these are not necessarily in common use and add both technical complexity and issues of unfamiliarity for users. Similarly, techniques exist to make it easier to trace values - for example Kennedy and Graham [48] switch straight polylines to curves (evaluated by Holten and van Wijk [49]). It is reasonable to assume that aspect ratio is likely to impact on some of these methods, although an evaluation of this is beyond the scope of this paper.
9 Supplemental materials
All supplemental materials are available on Figshare at https://dx.doi.org/10.6084/m9.figshare.31211107.
{acks}